OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x)= (x+1/2)(x+7)(x+7)(x+5)

jimthompson5910 (jim_thompson5910):

Hint: f(x)= (x+1/2)(x+7)(x+7)(x+5) 0 = (x+1/2)(x+7)(x+7)(x+5) (x+1/2)(x+7)(x+7)(x+5) = 0 x+1/2 = 0, x+7 = 0, or x+5 = 0 x = ??, x = ??, or x = ??